### 02 Apr 2026 v1

# When simulations look right but causal effects go wrong: Large language models as behavioral simulators

Abstract
Behavioral simulation is increasingly used to anticipate responses to interventions. Large language models (LLMs) enable researchers to specify population characteristics and intervention context in natural language, but it remains unclear to what extent LLMs can use these inputs to infer intervention effects. We evaluated three LLMs on 11 climate-psychology interventions using a dataset of 59,508 participants from 62 countries, and replicated the main analysis in two additional datasets (12 and 27 countries). LLMs reproduced observed patterns in attitudinal outcomes (e.g., climate beliefs and policy support) reasonably well, and prompting refinements improved this descriptive fit. However, descriptive fit did not reliably translate into causal fidelity (i.e., accurate estimates of intervention effects), and these two dimensions of accuracy followed different error structures. This descriptive-causal divergence held across the three datasets, but varied across intervention logics, with larger errors for interventions that depended on evoking internal experience than on directly conveying reasons or social cues. It was more pronounced for behavioral outcomes, where LLMs imposed stronger attitude-behavior coupling than in human data. Countries and population groups appearing well captured descriptively were not necessarily those with lower causal errors. Relying on descriptive fit alone may therefore create unwarranted confidence in simulation results, misleading conclusions about intervention effects and masking population disparities that matter for fairness.
Keywords Large language models $\cdot$ Behavioral simulation $\cdot$ Intervention effects $\cdot$ Climate psychology
Human response is one of the key unknowns when policies and interventions are designed [12, 30]. Behavioral simulation is widely used to estimate how attitudes and behaviors may change before those interventions are implemented. For decades, social scientists have turned their understanding of behavior into systems that can be used to reason about and simulate decision-making. They have done this through experiments and surveys [17, 26], through agent-based models [8, 27], and through statistical models of heterogeneity [33]. With these tools, researchers usually know what knowledge they have built into the system, and what rules the system follows in producing an inference [8]. However, with the emergence of large language models (LLMs), researchers can now describe a population and an intervention in natural language and instruct the model to infer how people might respond [3, 7, 9, 22]. The use of LLMs is expanding from text generation [5] and synthetic survey responses [32] into causal uses (e.g., experiment forecasting, message pre-testing, and policy simulation) [25, 4, 14]. How faithfully LLMs draw on and apply knowledge about human behavior, especially the part that governs intervention-induced change, becomes an unavoidable question.
Most existing evaluations of these systems stay at the descriptive level, e.g., comparing model-generated distributions [9], means [21], or group differences [28] with observed human data. Previous work has reported encouraging results [6, 2, 10], matching human data in political orientations [3] and climate attitudes [24]. Recent efforts to build dedicated behavioral simulators have demonstrated impressive descriptive fit across a range of cognitive tasks [6] and large-scale social dynamics [29]. In practice, such descriptive validation is often treated as evidence that a simulator is adequate for downstream uses that involve intervention forecasting or policy reasoning. However, a small but growing body of work that moves closer to causal evaluation do not sit comfortably with this assumption. Simulations that appear descriptively accurate have been found to systematically overestimate treatment effects [10, 18], recover only a fraction of observed experimental results even when individual-level prediction is strong [15, 36], or raise broader concerns about whether descriptive alignment alone is sufficient for downstream behavioral inference [20]. The resulting possibility is that descriptive fit and causal fidelity diverge. However, we still lack empirical evidence on when this divergence emerges, or whether it varies systematically with intervention logic, outcome domain, and population.
Figure 1: Roadmap of this study.
To address these questions, we take climate attitudes and behaviors as an informative test case, given their cross-national heterogeneity, the diversity of intervention logics, and the availability of large-scale cross-national experimental data. We use three LLMs from the GPT, Claude, and Gemini families, to simulate individual-level responses to 11 climate-psychology interventions for 59,508 respondents across 62 countries in the International Climate Psychology Collaboration (ICCP) dataset [11, 39], and we replicate the key analyses in two additional cross-national experimental datasets [34, 37]. For each respondent, we provide the model with demographic and socioeconomic characteristics, and instruct it to infer the climate beliefs, policy support, and pro-environmental action under each intervention. We further test whether prompting refinements that improve descriptive fit also improve the recovery of intervention effects. This design allows us to examine whether LLMs merely reproduce observed patterns (descriptive fit), or whether they use the information given to them in ways that support valid recovery of the direction and magnitude of intervention effects in randomized experiments (causal fidelity). Fig. 1 provides an overview.
Our analysis shows that LLMs can recover descriptive patterns in attitudes and cross-national differences reasonably well, and that prompting enhancements can further improve this descriptive fit. But matching observed response patterns does not reliably translate into accurate inference about how interventions change those responses. These distortions are systematic across outcome domains, intervention logics, prompting methods, and population groups. Our findings thus quantify a separation between descriptive fit and causal fidelity in LLM-based behavioral simulation, a pattern that held across three independent datasets. They also suggest that the information researchers provide to LLMs about respondents, context, and inferential goals does not consistently convert into reliable inference about behavioral change, underscoring the risks of treating descriptive fit as evidence of causal fidelity without direct causal evaluation. Targeted benchmarks that can show whether models are capturing causal effects are warranted.
Results
Descriptive fit in LLM simulations does not guarantee causal fidelity
In this study, we provided three models (GPT-4o-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku) with demographic and socioeconomic profiles of 59,508 respondents from 62 countries in the ICCP dataset. Each model was instructed to predict climate belief (0-100), policy support (0-100), and pro-environmental action (0-8, converted to 0-100). The same pipeline was applied to two additional cross-national experiments: Spampatti et al. [34], which covered 12 countries and 6 psychological inoculation interventions on climate disinformation, and Većkalov et al. [37], which contained 27 countries and 2 interventions on scientific-consensus communication.
Figure 2: A better descriptive fit does not necessarily imply better causal fidelity. a. Mean error (ME) by model and outcome under baseline conditions. b. Mean absolute error (MAE) by model and outcome. c. Country-level Spearman rank correlation and KNN overlap. d. Ratio of predicted to observed variance across models and outcomes. e. Enhancement selection pipeline, in which five candidate techniques were compared on a development sample (n = 5,093). KNN-based few-shot prompting and VBN-based chain-of-thought prompting were selected. f. Enhancement effect on country-level mean error. g. Enhancement effect on individual-level MAE. h. Enhancement effect on country-level Spearman correlation and KNN overlap. i. Enhancement effect on Jensen-Shannon divergence by outcome. j. Mean absolute ATE error by outcome and prompting method for each LLM.
At baseline, all three models recovered broad cross-national descriptive patterns, but performance varied strongly by outcome domain (Fig. 2a-d). Country-level mean error was modest for belief (0.24-0.65 percentage points, p.p.) and policy (-0.55 to 8.63 p.p.), but much larger for action, where GPT and Gemini overpredicted action by about 13.6 and 7.9 p.p., while Claude underpredicted it by 12.8 p.p. (Fig. 2a). Individual-level mean absolute error (MAE) showed similar patterns, remaining  16-18 points for belief and policy but rising to  40-43 points for action (Fig. 2b). Country-level structural alignment was also modest, with baseline Spearman correlations of 0.29-0.39 and KNN overlap of the most similar five countries of 0.21-0.25 (Fig. 2c). We also conducted baseline comparisons using multiple linear regression and LASSO (Supplementary Table S1), suggesting that LLMs could approach the individual-level descriptive performance of conventional statistical baselines despite operating without training examples.
Prompt enhancements further strengthened this descriptive fit. From a broader set of candidates spanning external information provision and internal reasoning elicitation (Fig. 2e; Supplementary Table S2), we selected two best-performing methods using the control group as a development sample (n = 5,093): KNN-based few-shot prompting (few-shot) and VBN-based chain-of-thought reasoning (VBN-CoT). Both improved descriptive fit. Country-level mean error decreased by 0.6-5.6 p.p. depending on outcome and method, and individual-level MAE gains were 0.5-1.4 p.p. (Fig. 2f-g). Few-shot produced larger gains in country-level alignment than VBN-CoT and raised Spearman correlation from about 0.39 to 0.65 (Fig. 2h). VBN-CoT produced larger reductions in distributional divergence for belief and policy, although neither method substantially improved distributional fit for action (Fig. 2i). At the same time, all models and prompting enhancements used exhibited variance compression, with predicted variance consistently below observed variance across all models and outcomes (Fig. 2d), indicating that descriptive fit was strongest for central tendencies rather than the full spread of human responses.
However, this descriptive fit did not carry over to causal fidelity. Under baseline prompting, when averaged across outcomes, mean absolute ATE error was 9.5 p.p. for GPT, 8.8 p.p. for Gemini, and 9.2 p.p. for Claude (Fig. 2j). These errors were especially larger for action, consistent with the descriptive results above. By contrast, linear regression and LASSO had lower mean absolute ATE errors of 4.3 p.p. and 4.2 p.p., respectively (Supplementary Table S1), suggesting that the limited causal fidelity of the LLM simulations was not solely a consequence of task difficulty.
Prompt refinements improved this picture selectively. VBN-CoT reduced mean absolute ATE error by 34.0%-41.2% relative to baseline across all three models, but these gains were concentrated in belief and policy (e.g., 12.5 p.p. and 6.5 p.p. to 3.9 p.p. and 3.5 p.p. for GPT), with little benefit for action (from 8.4 p.p. to 8.7 p.p.). Few-shot was highly model-dependent and in some cases worsened causal error for action (e.g., from 8.4 p.p. to 44.2 p.p. in Claude and from 16.4 p.p. to 36.7 p.p. in Gemini). In neither dataset did the prompting strategy that minimized individual-level MAE also minimize ATE error (Supplementary Table S3). Thus, the same prompting strategies that made the simulations look more descriptively reliable did not produce parallel improvements in causal fidelity.
Descriptive and causal errors are governed by different structures
Figure 3: Descriptive fit and causal fidelity reveal distinct error structures across models and datasets. a. Outcome type dominates descriptive error, but causal error redistributes across method, country, and their interactions. b. The structural shift replicates in both additional datasets, although the dimensions involved vary with experimental design. c. Intraclass correlation coefficients (ICC) show that country-level heterogeneity is amplified in causal evaluation across all model-dataset combinations.
The divergence between descriptive fit and causal fidelity also represents a difference in error structure (Fig. 3a). In descriptive evaluation, error variance was dominated by outcome type among all three models (93.3%-94.5%), and prompting method contributed less than 1%. However, the structure of causal error reorganized. The contribution of outcome type declined to 11.9%-40.8%, while method and outcome-by-method interactions jointly accounted for 24.1%-65.1% of variance. The three models showed similar qualitative shifts from outcome-dominated descriptive error to a more distributed causal error structure, although the dimensions most amplified differed somewhat across models. GPT showed greater country-related heterogeneity (49.9% from country and its interactions), Claude showed the strongest amplification of method and its interactions (65.1% combined), and Gemini retained both a relatively large outcome main effect (40.8%) and substantial method-related variance (40.2% combined for method and outcome-by-method interactions).
Similar structural shifts replicated in two additional datasets (Fig. 3b). Variance decomposition likewise showed that descriptive and causal error diverged along different dimensions in the two designs. Method played a larger role in causal than descriptive error (15.8% vs. 3.2%) in the Spampatti dataset, while in the Većkalov dataset the divergence was carried more by shifts in the relative contributions of outcome (from 2.7% to 28.4%) and country (from 64.1% to 49.7%). Across datasets, the divergence was robust, although the dimensions along which it emerged varied with experimental design.
Precision-weighted and multilevel analyses further specified this structural difference. Intraclass correlation coefficients (ICC) for country random intercepts increased from 0.8%-1.6% for MAE to 13.1%-16.1% for absolute ATE error (Fig. 3c), indicating substantially greater between-country heterogeneity in causal evaluations. ICCs were amplified in causal evaluations in both replication datasets (Spampatti from 0.3% to 29.2%, Većkalov from 4.9% to 53.9%). Permutation tests shuffling responses within country-by-outcome blocks further showed the structural difference between MAE and ATE error was unlikely to arise from block composition alone, in all three datasets (all p < 0.001). Therefore, descriptive fit and causal fidelity were shaped by different sources of error, so descriptive evaluation alone would miss some of the most consequential sources of causal error.
Figure 4: Causal error is structured by intervention logic, and LLM simulations impose stronger attitude-behavior coupling than observed in human. Results shown for GPT; other models in Supplementary Information. a. Situational simulation interventions show the largest causal overestimation, and cultural and group norms showed the smallest. Each point represents the signed ATE error of an intervention under baseline (red), few-shot (blue), or VBN-CoT (purple) prompting. Crosses indicate direction-flipped effects. Open circles indicate interventions whose human ATE 95% CI crosses zero. b. LLM simulations tighten the attitude-behavior coupling relative to human data. Lower triangles show individual-level response correlations. Upper triangles show intervention-level ATE correlations. c. Models are sensitive to persuasive force in intervention text for attitudinal outcomes, but action estimates do not respond coherently to the same perturbations. Bars show the change in absolute ATE error relative to the original intervention text under text perturbations. Error bars indicate 95% CI. d. Instructing models to decouple attitudes from behavior does not consistently reduce action ATE error. Corrections improve some interventions but overcorrect others.
Intervention-level causal error reveals differential sensitivity across outcomes
If the descriptive-causal divergence is structured, causal error would vary predictably with intervention logic. We examined this at the intervention level, focusing on GPT as the primary case and providing Gemini and Claude results in Supplementary Fig. S1 and  S2. Details of intervention logic and materials are provided in Supplementary Table LABEL:tab:s4.
The degree and stability of causal errors varied systematically across intervention categories (Fig. 4a). We inspected intervention-level signed ATE bias (deltadeltaATE = predicted ATE minus observed ATE) to distinguish overestimation from underestimation. Situational simulation interventions (e.g., future self-continuity, psychological distance) produced the largest overestimation for belief and policy (baseline mean deltadeltaATE: +10.05 and +6.24) and the most unstable action estimates, with several showing direction-flipped effects (crosses in Fig. 4a). Cultural and group norms interventions (e.g., dynamic social norm, system justification) exhibited the smallest bias overall (belief: +3.34; policy: +3.83; action: +0.64). Deductive reasoning interventions (collective action, scientific consensus, pluralistic ignorance) showed intermediate bias for belief and policy (baseline: +8.09 and +5.52), which VBN-CoT effectively reduced (+2.67 and +0.63). These differences indicate that causal error was structured jointly by intervention logic and outcome domain, with action exhibiting qualitatively different error patterns from belief and policy.
We also observed stronger couplings between attitudinal and behavioral outcomes in the LLM simulations than in the human data (Fig. 4b). For human responses, belief and action were only weakly coupled at both the individual level (r = 0.11) and the intervention level (ATE r = -0.03), and the policy-action correlation was similarly weak (ATE r = -0.14), consistent with the well-documented intention-behavior gap in environmental psychology. LLM simulations showed stronger cross-outcome correlations, especially between attitudes and behaviors. For instance, the belief-action response correlation increased to 0.39 under baseline, 0.49 under few-shot and 0.75 under VBN-CoT. For ATEs, belief-action ATE correlation increased from -0.03 in human data to 0.82 under VBN-CoT. Even though the three outcomes were elicited separately, LLM simulations preserved less separation across outcome domains than was observed in the underlying human data.
The contrast between attitudinal and behavioral outcomes became clearer when we changed intervention content and outcome framing separately (Fig. 4c-d). First, three representative interventions were edited to vary persuasive force while preserving overall structure: (1) surface paraphrase, (2) force removal, and (3) force reversal (Supplementary Tables LABEL:tab:s5a,  LABEL:tab:s5b, and  LABEL:tab:s5c). Force removal reduced the ATE overestimation for belief and policy, with the largest reductions for scientific consensus (belief: -4.49, policy: -3.68), intermediate reductions for dynamic norm (-2.32, -1.76), and the smallest reductions for future-self continuity (-1.15, -0.26). But ATE error for action changed less coherently under the same perturbations. Then, we altered how the attitude-behavior relation was framed without changing intervention content (Supplementary Table LABEL:tab:s5d). Instructing the model to infer attitudes and behaviors separately produced little change in belief or policy estimates and did not improve action ATE error consistently across interventions. Explicitly stating that behavioral change is typically smaller and slower than attitudinal change shifted action estimates, reducing ATE error for scientific consensus by 0.16 (from 1.36 to 1.20), leaving it largely unchanged for dynamic norm (-0.01, from 0.10 to 0.09), and overcorrecting for future-self continuity (+0.64, from 0.24 to 0.88, with bias flipping from overestimation to underestimation). These probes suggest that attitudinal simulations responded more consistently than behavioral ones to intervention content, and behavioral simulations were more sensitive to how the attitude-behavior coupling was framed. Therefore, causal distortion was especially marked when the simulation required linking attitudinal and behavioral domains.
Descriptive validation masks structured inequalities in causal fidelity
Figure 5: Countries and population groups well captured descriptively are not necessarily those with accurate causal estimates. Results shown for GPT; other models in Supplementary Information. a. Descriptive and causal error are weakly correlated at the country level. Open markers indicate baseline, and filled markers indicate prompting refinements. b. Countries with highest and lowest descriptive error. c. Countries with highest and lowest causal error. Overlap between b and c is minimal. d-e. Country-level descriptive gains concentrate in OECD and high-internet countries, but these advantages do not carry over to causal fidelity. Error bars show 95% CI. f-g. Descriptive disparities are largest along age and gender, while causal disparities are largest along political orientation and SES. Error bars show 95% CI.
Having shown that descriptive and causal fidelity do not share the same error structure, we next examined the practical consequences of this mismatch at the country and subgroup levels. We focus on GPT as the primary case in the main text. Claude and Gemini showed similar qualitative patterns, and their results are reported in the Supplementary Fig. S3 and S4.
At the country level, descriptive and causal error were only weakly associated (Fig. 5a). Baseline MAE and ATE error were only weakly associated among all three outcomes (belief r = 0.22; policy r = 0.12; action r = -0.08). This divergence was also visible in country rankings (Fig. 5b–c). The 10 lowest-error countries for MAE and ATE error shared only 2 members (Saudi Arabia and Venezuela), and the 10 highest-error countries overlapped on only 3 (Italy, China, and Portugal). In other words, a country that looked descriptively well captured was not necessarily one in which intervention effects were causally well recovered. Relying solely on descriptive evaluation would fail to identify the countries with the largest causal errors.
The same divergence changed which populations appeared most advantaged or most at risk under prompting improvements (Fig.  5d-g). At the country level, descriptive gains were modestly concentrated in OECD countries and countries with higher internet penetration, with between-group differences typically within 1-3 p.p. (e.g., VBN-CoT, OECD vs. non-OECD: 1.41 p.p. larger, p = 0.005; high vs. low internet: 1.99 p.p. greater, p = 0.002). On the causal side, these structures changed as between-group differences in ATE error were unstable and shifted direction across methods. OECD countries showed slightly higher causal gains under few-shot (difference: 0.68, p = 0.069) but lower gains under VBN-CoT (-0.75, p = 0.199). Country characteristics that appeared mildly advantageous at the descriptive level did not reliably translate into causal-level advantages. This divergence was more pronounced at the demographic level. Descriptive gains were largest along age and gender dimensions (e.g., 18-24 vs. 65+ age groups: 4.48 p.p. under few-shot and 5.28 p.p. under VBN-CoT; female groups consistently showed  1 p.p. greater MAE improvement than the male group; both p < 0.001). By contrast, the most prominent causal disparities appeared along political orientation (e.g. right-leaning vs. left-leaning: 7.74 p.p. under few-shot and 5.76 under VBN-CoT; p < 0.001) and SES (high vs. low: 3.55 and 2.92 p.p. for the two methods, p < 0.001), while gender and age differences were not significant. The inequality dimensions most salient in descriptive evaluation did not retain their prominence in causal fidelity.
Figure 6: Exploratory feature importance profiles for descriptive and causal error. Descriptive error changes are primarily associated with statistical prediction difficulty (e.g., true mean, original MAE), while causal error changes are more strongly linked to socioeconomic features (e.g., Gini coefficient, internet penetration). Signed SHAP percentage values from random forest models (100-seed averaged) predicting enhancement-induced changes in MAE and ATE error across outcomes. a. Few-shot. b. VBN-CoT. Positive values (red) indicate that higher feature values are associated with larger error increases; negative values (blue) indicate larger error reductions. Values were defined relative to the baseline prompt.
An exploratory SHAP analysis further indicated that enhancement-induced changes in descriptive and causal error were associated with different feature profiles (Fig. 6). Changes in MAE were primarily associated with statistical features. The normalized signed SHAP contribution of true mean to predicted changes in action MAE reached 74.6%, with original MAE and median age also contributing substantially. By contrast, changes in ATE error were more strongly associated with socioeconomic features. For instance, the SHAP value of Gini coefficient for belief ATE error was 31.6% (few-shot) and 25.0% (VBN-CoT), and the contributions of internet penetration rate and political orientation were also markedly elevated, although directions varied across outcomes and enhancement methods. Therefore, enhancement-induced changes in descriptive error were more closely linked to statistical prediction difficulty, whereas changes in causal error were more strongly associated with mismatches between national social structures and the model’s assumptions about how attitudes translate into behavior.
Discussion
This study examined how LLMs simulate human responses to behavioral interventions, using large-scale cross-national experimental data. The central question was how faithfully the information provided by researchers through population descriptions and intervention designs was taken up and used by the model. Our results show that LLMs can reproduce descriptive patterns in people’s attitudes and behaviors with reasonable accuracy, but are less reliable in simulating how these attitudes and behaviors changed under intervention. Our findings align with the possible divergence between descriptive fit and causal fidelity raised by earlier studies [10, 18, 15, 36, 20, 16], and go further by showing that this divergence has a systematic structure in outcome domain, intervention logic, and prompting method, and is consequential for how populations and interventions are judged. Descriptive error and causal error arise from different sources and are only weakly aligned across countries. Prompting strategies that improve descriptive fit can under some conditions increase causal bias. This divergence also varies across intervention types and alters the inequality patterns identified through descriptive criteria alone.
Changes in attitudes or behavioral intentions do not reliably translate into changes in actual behavior, and our results suggest that LLMs tend to tighten precisely this weak coupling. Both the training data and the learning objective may contribute to this tendency. Human text tends to present behaviors in narratively coherent and reason-giving forms [31], while the well-documented intention-behavior gap is not always explicitly represented in such accounts [38]. A model trained on such text is therefore exposed more often to cases in which attitudes and behavior appear aligned than to cases in which they come apart [19]. The next-token prediction objective that underlies current LLMs may further reinforce this asymmetry by favoring stable associations that recur throughout training corpora [13]. Intervention effects, however, depend on a different kind of knowledge (e.g., which specific populations change after a condition is introduced to what extent [1]). Such information appears less often in text in a stable or repeated form, making it harder for models to encode. Consequently, models may tend to treat attitude change as a more direct precursor of behavioral change than it is in practice, representing people as more coherent and consistent agents than they really are. VBN-CoT may further amplify this tendency by encouraging more explicit inferential linking across outcomes. Given prior findings that LLMs tend to overestimate treatment effects [10, 20], our results further suggest that the strongest overestimation may emerge when intervention effects depend on attitudes, beliefs, or norms being carried forward into behavior.
LLMs do not preserve all forms of intervention-relevant behavioral knowledge equally. They showed the smallest overall bias for cultural and group norms interventions, while showing larger overestimations for situational simulation interventions. Part of the explanation lies in the different ways these mechanisms are carried by text. For argument-based interventions, the persuasive work is largely done by the text itself (e.g., through the evidence and the logic laid out). For imagination- and simulation-based interventions, the text is closer to a prompt, and the driver of behavioral change is largely the mental experience it evokes, which differs from person to person and leaves much less of a trace in written language. Our probes, however, indicate that differences in how intervention mechanisms are represented in text do not explain everything. Prior work suggests that models often rely on surface matching at the descriptive level [28], but our causal results show that models are sensitive to persuasive cues in intervention texts. The difficulty begins later, when those cues have to be carried forward into simulated behavior. Bias in attitude-to-behavior translation could not always be corrected even when models were instructed that behavioral change is typically smaller and slower than attitudinal change. The uneven pattern across intervention types therefore appears to come from two sources. Some change mechanisms are more fully represented in text than others, and there is a consistent difficulty in calibrating the step from attitudinal to behavioral change.
Populations and societies that are well represented descriptively are not necessarily those whose responses to intervention are inferred more accurately. Our results show that the inequality structures at the descriptive and causal levels follow different patterns. Descriptive improvements were more concentrated along demographic patterns such as age and gender, while causal bias fell more heavily on political orientation and socioeconomic status. A similar mismatch appeared at the country level. Descriptive improvements were more concentrated in OECD countries and in countries with higher internet penetration, but these advantages did not carry over consistently to causal fidelity. Moreover, prompting strategies introduced to improve simulation quality also shift the distribution of error across groups and settings. Therefore, a group or society may be closely matched to real data in means and distributions and still have its response to intervention misread. Fairness here concerns both who is represented more accurately and whose mechanisms of change are better understood. If descriptive validation remains the primary standard, current fairness frameworks may not be sufficient to surface the real distributional risks of these systems.
As LLMs are increasingly used in multi-agent systems and larger-scale social simulations, the descriptive-causal divergence may spread along longer causal chains and become harder as it emerges into broader system dynamics. Clarifying how human knowledge about behavioral change is preserved, revised, or misused in LLMs is a precondition for judging whether such systems can be relied on in social decision making. We therefore suggest bringing independent causal checks into the design of studies that use LLM simulation to estimate intervention or policy effects. Different outcome domains and intervention logics also merit separate evaluation, since reasonable performance for attitudes does not guarantee comparable reliability for behavior. More targeted causal benchmarks that can show whether models are capturing causal effects and their heterogeneity are also warranted.
This work has several limitations. First, our core findings come from climate attitudes and behavior. Other domains may differ in the degree of attitude-behavior correlation and have different types of interventions. Whether the descriptive-causal divergence takes the same form in other behavioral domains warrants further testing. Second, we used lightweight versions of three LLMs and did not conduct fine-tuning. Future research is encouraged to explore whether this divergence exists in larger or newer models, as well as models fine-tuned for behavioral prediction. Third, although the ICCP dataset covers more than 60 countries, the national samples may not be uniformly representative of each country’s full population, so the human data used for causal fidelity may be more reliable in some settings than in others. Fourth, the causal fidelity assessed here is limited mainly to average treatment effects. Further work on recovering finer-grained causal quantities (e.g., heterogeneous effects or longer-term behavioral change) is warranted. Fifth, we mainly explore the potential mechanisms of the descriptive-causal divergence through external response patterns and do not observe them directly in the models’ internal representations. Future work could introduce more fine-grained probing and examine model parameters and training processes to clarify the mechanisms behind the divergence more explicitly.
Materials and Methods
Data
We used the International Climate Psychology Collaboration (ICCP) dataset, a pre-registered cross-national randomized experiment in which 59,508 participants from 62 countries were randomly assigned to one of 11 behavioral interventions or a no-intervention control [11, 39]. We retained three continuous outcomes for analysis: 1) climate change belief (4 items on a 0-100 scale), 2) policy support (9 items on a 0-100 scale), and 3) pro-environmental action (via the Work for Environmental Protection Task [23], 0-8 trees planted and standardized to 0-100).
We replicated core findings on two additional climate-related experimental datasets: Spampatti et al. [34] and Većkalov et al. [37]. We retained outcomes that most closely mapped onto the focal domains of the ICCP analysis. In the Spampatti dataset, we used climate belief (9 items on a 1-5 scale) and the same Work for Environmental Protection Task as belief and action outcomes, respectively, all rescaled to 0-100. In the Većkalov dataset, we used perceived scientific consensus (2 items on a 0-100 scale), belief and concern (4 items on a 1-7 scale), and policy support (1 item on a 1-7 scale), all rescaled to 0-100. Full variable lists are provided in Supplementary Information.
LLM simulation pipeline
Models and prompting. Based on what the survey included, we selected four sets of variables as background information for LLMs: 1) Demographic information, including age, gender, country, and education; 2) Socio-economic status, including a direct material-asset measure based on household appliance ownership and an indirect subjective-status measure based on the MacArthur ladder; 3) Political orientation on social issues and economic issues; and 4) Perceived scientific consensus regarding human-caused climate change. The background information was serialized into narrative prompts to instruct LLMs. For instance:
“#YOUR PROFILE: A 39-year-old man living in the US with an education of college/university education (13-16 years) where the person answered the question ’What is your political orientation for social issues? (0-Extremely liberal/left-wing to 100-Extremely conservative/right-wing)’ as ’87’, the person indicated they own/have access to: Separate room for kitchen, Washing machine, Vacuum cleaner, Freezer/deep freeze, Personal computer, Bathroom, and Television, the person answered the question ’To the best of your knowledge, what percentage of climate scientists have concluded that human-caused climate change is happening? (0-100 Percentage)’ as ’51’……”
For participants in intervention conditions, the prompt additionally included the intervention text as presented to the original survey respondent. The 11 interventions span three theoretical categories: 1) deductive reasoning, including collective action framing, scientific consensus communication, pluralistic ignorance correction; 2) situational simulation techniques, including writing a letter to future generation, writing a letter as the future yourself, psychological distance reduction, and negative emotion induction; and 3) cultural and group norms, including system justification reframing, building moral foundations, dynamic social norms, working-together norms. The explanations of the interventions are provided in Supplementary Table LABEL:tab:s4, and the psychological rationales of designing these interventions could be found at Vlasceanu et. al [39]. The converted intervention texts in the prompts are provided in Supplementary Information.
These prompts were provided to three widely-used LLMs in large-scale simulations, including GPT-4o-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku through their APIs. Models were instructed to return item-level responses in JSON format. We ran separate simulations for the three outcome domains. All models were queried with temperature = 0.
Prompting refinements. We developed and tested two categories of enhancement strategies on a development set drawn from the control group (n = 5,093). External information methods augmented prompts with real-world data, including KNN-based few-shot retrieval (demographically similar exemplars within the sample’s country), country-level distributional context (socio-economic, demographic, and climate-related information), and target-variable distributional statistics (national means, variances, and median). Internal reasoning methods guided the model’s inference process: generic chain-of-thought and Value-Belief-Norm (VBN) theory-based chain-of-thought. Details of these methods and complete prompt templates are provided in Supplementary Text S1-S3.
Based on development set performance (MAE), we selected two enhancement methods for full evaluation as illustrated in Fig. 2e: 1) Few-shot: the prompt was augmented with 6 demographically similar exemplars retrieved via KNN within the same country, experimental condition, and outcome domain as the target case. Each exemplar included its observed response on the corresponding outcome, and the target respondent was excluded from retrieval. 2) VBN-based chain of thought (VBN-CoT): the model first inferred the profile’s environmental values, climate change beliefs, and personal norms based on their demographic profile (stage 1), then predicted outcome responses conditioned on this intermediate reasoning (stage 2). This design operationalizes the Value-Belief-Norm theory of environmentalism [35], which posits that pro-environmental behavior arises from a causal chain of values – ecological worldview – awareness of consequences – personal norms – behavior. The baseline condition included the full serialized participant profile, without additional retrieved exemplars or reasoning guidance.
The same prompting templates and enhancement methods were applied to the replication datasets.
Model outputs. Each participant was simulated under all three methods. For each simulated participant, we recorded the predicted response on each outcome. Control and treatment participants were simulated using the same profile template. For treatment participants, the assigned intervention text was appended to the prompt.
Evaluation
We evaluated simulation accuracy along three dimensions to capture different aspects of fidelity.
Country-level descriptive fit. Mean error (ME), defined as the signed difference between predicted and observed country-level means for each outcome and method. ME captures systematic directional bias, i.e., whether the model consistently over- or underestimates a country’s average attitude. To provide a broader picture of model performance across contexts, we also computed country-level: 1) Spearman rank correlations, which measure whether the model preserves the relative ordering of countries; 2) KNN overlap (k = 5), which quantifies whether the model preserved country similarity structure. For each country, we constructed a country profile from item-level distributional summaries, identified its k nearest countries in the observed country-profile space and in the model-predicted country-profile space, and computed the overlap between the two neighbor sets; and 3) Jensen-Shannon divergence, which assesses distributional similarity between predicted and observed responses within each country. It was calculated at the country-by-item level from task-specific binned response distributions and then aggregated across items.
Individual-level descriptive fit. For each participant and construct, we first averaged item-level predictions and observed responses within the construct, then computed the absolute difference between the predicted and observed construct scores. MAE was defined as the mean of these person-level absolute errors within each cell defined by country, outcome, and method. MAE captures how well the model reproduces individual-level variation in attitudes.
Causal fidelity. For each country, outcome, method, and intervention cell, we computed predicted ATEs as the mean difference between simulated treatment and control groups, and compared these against ground-truth ATEs from the ICCP randomized experiment. ATE error was defined as the absolute difference between predicted and ground-truth ATEs (|predicted ATE – true ATE|), capturing the magnitude of causal estimation error regardless of direction. We use this unsigned metric throughout the main text and figures unless otherwise noted. Standard errors of ATE error estimates were computed for precision-weighted analyses (see Statistical analysis).
Conventional statistical models. We compared LLM performance on both descriptive fit and causal fidelity against conventional statistical approaches using MAE and absolute ATE error. We fitted linear regression and LASSO ($\alpha$ = 1.0) models for each outcome domain using the same demographic and socioeconomic features provided to the LLMs. Models were trained separately for each experimental condition (including control). Within each condition, data were split 80/20 into training and test sets. We used 5-fold cross-validation on the training split to estimate out-of-sample performance.
Statistical analysis
All analyses compare two dependent variables: individual-level MAE (descriptive fit) and absolute ATE error (causal fidelity). Because the replication datasets had more limited designs, including fewer countries, interventions, and outcome domains, we restricted replication analyses to variance decomposition, mixed-effects ICC estimation, and permutation tests.
Variance decomposition. We decomposed the variance in MAE and ATE error using factorial ANOVA. For MAE, individual-level data were aggregated to country-outcome-method cells and modeled with country, outcome, method, and all pairwise interactions as fixed effects. For ATE error, we conducted two analyses: 1) a comparable country-aggregated model using the same factor structure, and 2) an intervention-level model additionally including intervention, method-by-intervention, and outcome-by-intervention terms. All models used Type II sums of squares, with $\eta^{2}$ factor sum of squares divided by total sum of squares. Here we chose Type II because it tests each factor after accounting for all other factors at the same or lower order, which is appropriate given our unbalanced design. Type III sums of squares were computed as a robustness check and yielded consistent conclusions.
Mixed-effects models. To account for the clustering by countries, we fitted linear mixed-effects models with country as a random intercept using REML estimation. For MAE, the model included outcome, method, and their interaction as fixed effects. For ATE error, fixed effects included outcome, method, intervention, and all pairwise interactions. We report variance components and intraclass correlation coefficients (ICC, defined as country variance divided by total variance) to quantify the contribution of country-level heterogeneity to each metric.
Permutation test. To validate the method-related variance component without relying on standard distributional assumptions, we conducted a permutation test with 5,000 iterations. In each iteration, method labels were shuffled within country-outcome blocks, the full factorial ANOVA was refitted, and the sum of all method-related sum of squares terms (method main effect, outcome-by-method, and country-by-method) was extracted as the test statistic. The permutation p-value was the proportion of permuted method block sum-of-squares values greater than or equal to the observed value.
Descriptive-causal correlation. To assess whether descriptive fit serves as a proxy for causal accuracy, we computed Spearman correlations between country-level MAE and ATE error within each outcome-by-method combination. Country-level MAE was defined as the mean individual MAE within a given country, and each correlation was computed across the 62 countries.
Cross-task correlation. To quantify whether LLM simulations impose stronger cross-outcome coherence than observed in human data, we computed Pearson correlation matrices among belief, policy, and action using individual-level construct scores for human responses and for each simulation method. We also computed correlation matrices at the ATE level across countries, correlating country-level ATEs across the three outcomes.
Perturbation analysis
To probe whether LLM causal fidelity reflects sensitivity to intervention content versus artifacts of attitude-behavior coupling, we conducted two sets of perturbation experiments on three representative interventions, one from each theoretical category: scientific consensus (deductive reasoning), dynamic social norm (cultural and group norms), and future-self continuity (situational simulation). All perturbation analyses were conducted using GPT.
In the first set, we modified intervention texts along a gradient of persuasive force while preserving overall structure. For each intervention, three perturbed versions were generated using GPT-5.4 thinking with explicit editing instructions: (1) surface paraphrase, in which the original text was reworded without altering argumentative content or persuasive force; (2) force removal, in which persuasive elements (e.g., specific evidence, emotional appeals, or normative cues) were removed while retaining the topic and framing; and (3) force reversal, in which the direction of the persuasive argument was reversed (e.g., undermining rather than supporting pro-environmental action). Generated texts were reviewed by the authors to ensure that only the intended dimension was altered. For each perturbation, the full simulation pipeline was re-run under baseline prompting, and the change in absolute ATE error relative to the original intervention text was computed for each outcome domain.
In the second set, we held intervention content constant and varied how the model was instructed to relate attitudinal and behavioral outcomes. Two framing conditions were tested: (1) separate inference, in which the model was explicitly instructed to infer attitudes and behaviors independently; and (2) calibrated behavior, in which the prompt included an explicit statement that behavioral change is typically smaller and slower than attitudinal change. Changes in absolute ATE error relative to the unmodified prompt were computed for each outcome domain. The complete original and perturbed intervention texts are provided in Supplementary Information.
Inequality and feature importance analysis
Country-level inequality. To assess whether cross-national reliability gaps change when moving from descriptive to causal fidelity, we stratified countries by four structural characteristics: OECD membership (binary, to capture economic development), internet penetration (tercile split, to capture digital access), English proficiency (tercile split, to capture potential linguistic alignment with model training data), and income inequality (Gini coefficient, tercile split, to capture socio-economic conditions). For each stratification, between-group comparisons focused on the extreme groups (e.g., High vs. Low for English proficiency, top vs. bottom tercile for internet penetration and Gini). We compared between-group differences in mean MAE / mean ATE error and estimated 95% bootstrap confidence intervals based on 5,000 resamples. We additionally used Welch’s t-test and Mann-Whitney U test to assess the robustness of group differences. Except for English proficiency scores, which were obtained from the EF English Proficiency Index (EF EPI), all other country-level data were collected from Our World in Data.
Demographic subgroup analysis. To assess whether descriptive fit and causal fidelity identify the same groups as relatively well or poorly simulated, we examined simulation accuracy across four demographic contrasts: gender (male vs. female), subjective socio-economic status (high vs. low), political orientation (left vs. right), and age (18–24 vs. 65+). These contrasts were chosen to represent salient axes of social and political heterogeneity that are also substantively relevant in climate-attitude research. For each contrast, we computed mean MAE and mean ATE error and compared the size of between-group disparities across descriptive fit and causal fidelity using the same bootstrap and significance-testing procedures as in the country-level inequality analysis.
Feature importance exploration. As an exploratory extension, we trained separate Random Forest models to predict enhancement-induced changes in individual-level MAE and country-level ATE error relative to baseline, and used SHAP values to characterize feature-level correlates of descriptive and causal reliability disparities. Models were fit at the country-by-outcome-by-method level using country-level characteristics (internet penetration, English proficiency, Gini coefficient), demographic composition (median age, gender ratio, proportion high-SES, proportion left-leaning), and statistical properties of the outcome (true mean, true standard deviation, original MAE) as predictors. Thus, the SHAP analysis incorporated all grouping variables used in the inequality analyses, along with additional statistical properties of each outcome. We then computed SHAP values to quantify each feature’s directional contribution to the predicted change in error relative to baseline, averaging across 100 random seeds to stabilize estimates.
Acknowledgements
FJ was supported by the Connaught Fund (Grant Number 520245) and Social Sciences and Humanities Research Council (SSHRC) of Canada (Grant Number 215119, 00169), Seed Grant of the International Network of Educational Institutes (Grant Number 522011) and Canada Research Chair program (CRC-2024-00169). The authors would like to thank Zhixun Yan and Yuxuan Yang from the University of Toronto for their research assistance.
Conflict of interest
The authors declare no competing interests.
Supplementary Information
Table S1: Individual-level MAE and absolute ATE error of conventional statistical models and LLMs.
The results were pooled for the 12 conditions (control and 11 intervention groups). The LLMs operated in a zero-shot setting without access to training examples, whereas OLS and LASSO were supervised baselines trained on an 80/20 split and evaluated on held-out test data. Because the statistical baselines were evaluated on held-out test sets and sample sizes varied across countries and conditions, country-level comparisons were less directly comparable across methods. We therefore focus this comparison on individual-level metrics.
Metric OLS LASSO LLM (GPT) LLM (Gemini) LLM (Claude)
Individual MAE 22.98 23.65 24.24 24.53 26.10
   Belief 15.26 15.40 15.99 16.22 17.96
   Policy 13.92 14.05 16.11 15.23 17.22
   Action 39.75 41.50 40.63 42.13 43.13
Absolute ATE error 4.32 4.24 9.47 8.85 9.23
   Belief 2.99 2.97 7.83 6.52 12.53
   Policy 2.87 2.85 5.93 3.42 6.52
   Action 7.10 6.90 14.65 16.59 8.64
Table S2: Development-set performance of five enhancement candidates used for method selection ($n$ = 5,093).
Few-shot supplied responses from demographically similar respondents retrieved within the same country and experimental condition using $K$-nearest neighbor clustering ($k$ = 6). VBN-CoT used a two-stage procedure in which the model first inferred a Value-Belief-Norm profile and then generated survey responses conditioned on that intermediate plan. Country-context and distributional-statistics prompts appended country-level contextual summaries or target-variable distributional benchmarks, respectively. Generic chain-of-thought prompting used a two-stage reasoning plan without VBN-specific structure. Based on development-set performance in individual-level MAE, we selected few-shot and VBN-CoT for full evaluation in the main analyses. Complete implementation details and exact prompt templates are available in the analysis code.
External information Internal inference
Metric Baseline Few-shot Country context Distributional stats CoT VBN-CoT
Individual MAE 23.82 22.86 24.09 23.67 24.91 23.77
   Belief 15.79 15.80 15.59 15.90 15.59 15.68
   Policy 15.41 14.84 15.19 14.56 16.41 16.41
   Action 40.25 37.92 41.48 40.56 42.72 39.21
Table S3: Prompting-method comparison in the Spampatti and Većkalov replication datasets.
Dataset Method Individual MAE MAE rank Absolute ATE error ATE-error rank
Spampatti Baseline 28.07 2 26.52 3
Spampatti Few-shot 21.09 1 19.28 2
Spampatti VBN-CoT 29.70 3 4.30 1
Većkalov Baseline 10.61 2 2.04 2
Većkalov Few-shot 10.60 1 2.07 3
Većkalov VBN-CoT 12.01 3 2.00 1
Table S4: Overview of the 11 climate-psychology interventions used in the ICCP simulations. Adapted from Figure 1 in Vlasceanu et al. (2024).
Category Intervention Description
Deductive reasoning Collective action Features examples of successful collective action that have had meaningful effects on climate policies (e.g., protests) or have solved past global issues (e.g., the restoration of the ozone layer).
Scientific consensus Informs participants that “99% of expert climate scientists agree that Earth is warming and climate change is happening, mainly because of human activity.”
Pluralistic ignorance Presents real public opinion data collected by the United Nations that show what percentage of people in each participant’s country agree that climate change is a global emergency.
Situational simulation Letter to future generations Emphasizes how one’s current actions affect future generations by asking participants to write a letter to a socially close child who will read it in 25 years when they are an adult, describing current actions toward ensuring a habitable planet.
Future self-continuity Emphasizes the future self-continuity by asking each participant to project themselves into the future and write a letter addressed to themselves in the present, describing the actions they would have wanted to take regarding climate change.
Psychological distance Frames climate change as a proximal risk using examples of recent natural disasters caused by climate change in each participant’s nation and prompts them to write about the climate impacts on their community.
Negative emotions Exposes participants to ecologically valid scientific facts regarding the impacts of climate change framed in a “doom and gloom” style of messaging that were drawn from different real-world news and media sources.
Cultural & group norms System justification Frames climate change as threatening to the way of life to each participant’s nation and makes an appeal to climate action as the patriotic response.
Building moral foundations Invokes authority (e.g., “from scientists to experts in the military, there is near universal agreement”), purity (e.g., keep our air, water, and land pure), and ingroup-loyalty (e.g., “it is the American solution”) moral foundations.
Dynamic social norm Informs participants of how country-level norms are changing and “more and more people are becoming concerned about climate change,” suggesting that people should take action.
Work-together norm Combines referencing a social norm (i.e., “a majority of people are taking steps to reduce their carbon footprint”) with an invitation to “join in” and work together with fellow citizens toward this common goal.
Table S5: Perturbed texts for the Scientific Consensus intervention. Underlined bold text indicates key modified phrases.
Original Surface paraphrase Force removal Force reverse
Did you know that 99% of expert climate scientists agree that the Earth is warming and climate change is happening, mainly because of human activity (for example, burning fossil fuels)? Did you know that 99% of expert climate scientists agree that global warming is accelerating and the climate is changing, primarily driven by human actions (for example, fossil fuel combustion)? Did you know that 99% of expert climate scientists agree that the Earth is warming and climate change is happening, while its causes are still under study (for example, temperature records)? Did you know that 99% of expert climate scientists agree that the Earth is warming and climate change is happening, mainly because of natural cyclical processes (for example, solar cycles)?
Table S6: Perturbed texts for the Dynamic Social Norm intervention. Omitted passages are identical to the original. Underlined bold text indicates key modified phrases.
Original Surface paraphrase Force removal Force reverse
People in {country} and around the world are changing: more and more people are concerned about climate change, and are now taking action across multiple fronts. Since 2013, concerns about climate change have increased in most countries surveyed. People in {country} and around the world are shifting: a growing number of people are worried about climate change, and are now acting on multiple fronts. Since 2013, worries about climate change have risen in most countries surveyed. People in {country} and around the world hold a range of views on climate change, and show different patterns of engagement across multiple fronts. Since 2013, views on climate change have been recorded across most countries surveyed. People in {country} and around the world are changing: fewer and fewer people are concerned about climate change, and fewer are taking action across multiple fronts. Since 2013, concerns about climate change have declined in most countries surveyed.
[Chart: Pew Research Center, Spring 2013 & 2018 Global Attitudes Surveys. For each country, paired data points connected by a line show that concern has grown over the period.] [Chart: Pew Research Center, Spring 2013 & 2018 Global Attitudes Surveys. For each country, paired data points connected by a line show that worry has risen over the period.] [Chart: Pew Research Center, Spring 2013 & 2018 Global Attitudes Surveys. For each country, paired data points connected by a line show how views have been tracked over the period.] [Chart: Pew Research Center, Spring 2013 & 2018 Global Attitudes Surveys. For each country, paired data points connected by a line show that concern has declined over the period.]
What kinds of actions are people taking right now? More than ever before, people are making changes to their lifestyles, supporting policies to address climate change, and are giving the issue more time and attention. For example, more and more people from around the world are now: [Practice unchanged] What kinds of actions are people taking right now? More than before, people are adjusting their lifestyles, backing policies to address climate change, and are devoting more time and attention to the issue. For example, a rising number of people from around the world are now: [Practice unchanged] What kinds of patterns can be observed? People show different patterns in their lifestyles, views on policies related to climate change, and the time and attention they give to the issue. For example, people from around the world show varied patterns in: [Practice unchanged] What kinds of actions are people taking right now? Compared with before, fewer people are making changes to their lifestyles, supporting policies to address climate change, and giving the issue time and attention. For example, fewer and fewer people from around the world are now: [Practice unchanged]
There’s also been a notable increase in support for climate change mitigation policy–some of the most popular policies include: [Policies unchanged] There’s also been a notable increase in support for climate change mitigation policy–some of the most popular policies include: [Policies unchanged] Views on climate change mitigation policy also vary across contexts–some of the most discussed policies include: [Policies unchanged] There’s also been a notable decline in support for climate change mitigation policy–some of the least popular policies include: [Policies unchanged]
Table S7: Perturbed texts for the Future-Self Continuity intervention. Omitted passages are identical to the original. Underlined bold text indicates key modified phrases.
Original Surface paraphrase Force removal Force reverse
[Climate change description unchanged] [Climate change description unchanged] [Climate change description unchanged] [Climate change description unchanged]
[Climate and human activity unchanged] [Climate and human activity unchanged] [Climate and human activity unchanged] [Climate and human activity unchanged]
Please put yourself in the year 2030 - eight years from now. Take a few moments to imagine your life in that future. Imagine how you will look, where you will be, and who you are with. In the year 2030, it will be clear whether keeping climate change under 2C is still possible. It will be clear whether the necessary change occurred fast enough to match the speed of the changing climate. As the Earth’s atmosphere continues to heat up, the effects of climate change will be more apparent: the “highest observed temperature” records will keep being updated, heatwaves and the draughts will become more common, species will continue to become extinct. Now picture yourself in the year 2030 - eight years into the future. Take a few moments to envision your life at that point in time. Think about how you will look, where you will be living, and who will be around you. By 2030, it will be evident whether limiting warming to 2C remains an achievable goal. It will be clear whether the necessary changes happened quickly enough to keep pace with the changing climate. As temperatures continue to climb, the effects of climate change will become more apparent: temperature records will keep being broken, heatwaves and droughts will intensify and spread, and more species will face the threat of extinction. Please put yourself in the year 2030 - eight years from now. Take a few moments to imagine your life in that future. Imagine how you will look, where you will be, and who you are with. In the year 2030, it will be clearer how likely keeping climate change under 2C still appears to be. It will be clear whether various factors played out as projected in current climate models. As the Earth’s atmosphere continues to heat up, the long-term effects of climate change will become clearer over time: the “highest observed temperature” records will keep being updated, heatwaves and the droughts will become more common, species will continue to become extinct. Please put yourself in the year 2030 - eight years from now. Take a few moments to imagine your life in that future. Imagine how you will look, where you will be, and who you are with. In the year 2030, the trajectory of climate change will have been shaped by forces beyond any individual’s control. It will be clear whether the changing climate was shaped mainly by forces beyond individual action. As the Earth’s atmosphere continues to heat up, the effects of climate change will unfold regardless of individual choices: the “highest observed temperature” records will keep being updated, heatwaves and the droughts will become more common, species will continue to become extinct.
Table S8: Prompt framings used in the attitude-behavior coupling probe. Underlined bold text indicates the inserted framing statement.
Original Coupling removal Coupling attenuation
Here is your profile:
{persona}
{intervention}
{question}
Here is your profile:
{persona}
{intervention}
Please note, people’s views on climate change and their everyday behaviors are related but distinct, and should be inferred separately.
{question}
Here is your profile:
{persona}
{intervention}
Please note, even when people’s views on climate change shift, changes in everyday behavior are often smaller/slower because of habit/cost/social context.
{question}
Table S9: Overview of participant profile variables (prompt inputs) and outcome variables across datasets.
All outcomes were rescaled to a common 0–100 metric for cross-study comparison.
Dataset Prompt input variables Outcome variables
ICCP
(62 countries,
11 interventions)
Age, gender, country, education, socioeconomic status (direct material-asset measure and indirect subjective-status measure), political orientation (social and economic issues), perceived scientific consensus Climate belief (4 items, 0–100)
Policy support (9 items, 0–100)
Pro-environmental action (WEPT, 0–8 $\rightarrow$ 0–100)
Spampatti et al.
(12 countries,
6 interventions)
Age, gender, country, education, political ideology, cognitive reflection (CRT-2) Climate belief (9 items, 1–5 $\rightarrow$ 0–100)
Pro-environmental action (WEPT, 0–8 $\rightarrow$ 0–100)
Većkalov et al.
(27 countries,
2 interventions)
Age, gender, country, urbanicity, student status, college-degree group, political ideology, trust in climate scientists Perceived consensus (2 items, 0–100)
Belief and concern (4 items, 1–7 $\rightarrow$ 0–100)
Policy support (1 item, 1–7 $\rightarrow$ 0–100)
Table S10: Results of Welch’s $t$-test for group comparisons (MAE). ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$ when both tests were significant; indicates that only one of the two tests reached $p<0.05$; unmarked cells were not significant in either test.
GPT Gemini Claude
Group Contrast Few-shot VBN-CoT Few-shot VBN-CoT Few-shot VBN-CoT
Gini High vs. Low 0.76-0.76 +0.48+0.48 $4.91^{**}$ 0.58-0.58 $2.96^{*}$ +0.76+0.76
Internet rate High vs. Low +0.57+0.57 0.92-0.92 $3.37^{*} +$$1.51^{*} +$1.72+1.72 $1.28^{*}$
English proficiency High vs. Low 0.15-0.15 0.39-0.39 $5.13^{**} +$$1.97^{**} +$1.55+1.55 $1.02^{\dagger}$
OECD OECD vs. Non-OECD +0.68+0.68 0.75-0.75 $3.39^{**} +$$26^{*} +1.1$9+1.19 $1.20^{**}$
Age 18–24 vs. 65+ +1.06+1.06 +0.97+0.97 $4.84^{***}$ $0.91^{\dagger}$ $2.44^{*} +$0.14+0.14
Gender Male vs. Female 0.12-0.12 0.36-0.36 +0.45+0.45 +0.15+0.15 $1.22^{*} +$0.24+0.24
Political orientation Left vs. Right $7.74^{***}$ $5.76^{***}$ $2.84^{**} +$0.27+0.27 $3.89^{***}$ $5.48^{***}$
SES High vs. Low $3.55^{***} +$$2.92^{***}$ $0.88^{\dagger}$ $1.05^{*}$ $1.66^{*}$ 0.02-0.02
Table S11: Results of Welch’s $t$-test for group comparisons (ATE). ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$ when both tests were significant; indicates that only one of the two tests reached $p<0.05$; unmarked cells were not significant in either test.
GPT Gemini Claude
Group Contrast Few-shot VBN-CoT Few-shot VBN-CoT Few-shot VBN-CoT
Gini High vs. Low 0.05-0.05 0.15-0.15 0.21-0.21 $2.55^{**}$ 0.10-0.10 $2.32^{***}$
Internet rate High vs. Low $1.25^{*} +$$1.99^{**} +$0.06+0.06 +1.24+1.24 +1.01+1.01 $2.66^{***}$
English proficiency High vs. Low +0.27+0.27 +1.09+1.09 +0.13+0.13 $2.14^{*} +$0.27+0.27 $2.42^{***}$
OECD OECD vs. Non-OECD +0.58+0.58 $1.41^{**}$ 0.16-0.16 $1.32^{*} +$0.62+0.62 $2.20^{***}$
Age 18–24 vs. 65+ $4.48^{***}$ $5.28^{***}$ $2.56^{***}$ $6.67^{***}$ $4.98^{***}$ $0.57^{\dagger}$
Gender Male vs. Female $0.98^{***}$ $1.10^{***}$ $0.66^{*}$ $1.63^{***}$ $1.01^{***}$ $0.62^{**}$
Political orientation Left vs. Right $1.90^{***}$ +0.11+0.11 $0.97^{*} +$$2.09^{***} +$0.36+0.36 $1.58^{***}$
SES High vs. Low +0.06+0.06 +0.41+0.41 $2^{\dagger} +0.77+0$.77 $0.78^{*}$ $0.92^{**}$
Supplementary Text S1 : Baseline prompt template.
System You are a survey participant. You will be given a profile that contains factual background and prior questionnaire responses, as well as materials that are part of the survey context. Your goal is to complete the survey in a way that reflects how this specific person would likely think, feel, and respond in this situation, given their background and the materials they are presented with. Prioritize consistency with the person’s background and prior responses rather than general norms or idealized answers. Follow the response format EXACTLY as requested (e.g., JSON only). No extra keys, no comments, no markdown, no trailing text. Treat each question as part of the same survey session; keep answers internally consistent with the profile and your previous responses.
User (belief as an example) # YOUR PROFILE
[PROFILE]
# SURVEY QUESTION
How accurate do you think these statements are? Answer on a scale from 0 to 100, where 0 means “not at all accurate” and 100 means “extremely accurate”.
Questions: Q1: Taking action to fight climate change is necessary to avoid a global catastrophe. Q2: Human activities are causing climate change. Q3: Climate change poses a serious threat to humanity. Q4: Climate change is a global emergency. # RESPONSE FORMAT
Respond in the following JSON format only, with no additional text:
{"Q1": <number>, "Q2": <number>, "Q3": <number>, "Q4": <number>}
Supplementary Text S2 : Few-shot prompt template
System Same as baseline.
User (belief as an example) # YOUR PROFILE
[PROFILE]
# SURVEY QUESTION
How accurate do you think these statements are? Answer on a scale from 0 to 100, where 0 means “not at all accurate” and 100 means “extremely accurate”.
Questions: Q1: Taking action to fight climate change is necessary to avoid a global catastrophe. Q2: Human activities are causing climate change. Q3: Climate change poses a serious threat to humanity. Q4: Climate change is a global emergency. # RESPONSE FORMAT
Respond in the following JSON format only, with no additional text:
{"Q1": <number>, "Q2": <number>, "Q3": <number>, "Q4": <number>}
# REFERENCE
Reference answers from similar participants (same country & condition):
– Example 1: [FEATURE_1=value | FEATURE_2=value | FEATURE_3=value] -> {"Q1": ..., "Q2": ..., "Q3": ..., "Q4": ...}
– Example 2: [FEATURE_1=value | FEATURE_2=value | FEATURE_3=value] -> {"Q1": ..., "Q2": ..., "Q3": ..., "Q4": ...}
– Example 6: [FEATURE_1=value | FEATURE_2=value | FEATURE_3=value] -> {"Q1": ..., "Q2": ..., "Q3": ..., "Q4": ...}
Supplementary Text S3 : VBN-CoT prompt template
Stage 1 – System You are an expert of the Value-Belief-Norm (VBN) framework for pro-environmental behavior. You are preparing to answer a survey question as a specific person. You will be given a profile that contains factual background and prior questionnaire responses. Your goal is to infer the person’s value orientation (Schwartz self-enhancement $\leftrightarrow$ self-transcendence) and then extract VBN components (Awareness of Consequences – AC; Ascription of Responsibility – AR; Personal Norm – PN) from this profile. Output JSON only. No extra keys, no comments, no markdown, no trailing text.
Stage 1 – User (belief as an example) # Profile:
[PROFILE]
# Question:
How accurate do you think these statements are? Answer on a scale from 0 to 100, where 0 means “not at all accurate” and 100 means “extremely accurate”.
Questions: Q1: Taking action to fight climate change is necessary to avoid a global catastrophe. Q2: Human activities are causing climate change. Q3: Climate change poses a serious threat to humanity. Q4: Climate change is a global emergency. Respond in the following JSON format only, with no additional text: {"Q1": <number>, "Q2": <number>, "Q3": <number>, "Q4": <number>}
Return a VBN plan based on the profile for judging the accuracy of climate-change statements:
1) evidence: list 2–4 most relevant explicit profile items (must be explicit; no guessing). 2) values: Based on this person’s political orientation, social position, and education, infer their dominant value orientation on the Schwartz self-enhancement $\leftrightarrow$ self-transcendence dimension. – self-enhancement: prioritizes personal status, wealth, and authority – self-transcendence: prioritizes welfare of others, nature, and equality – mixed: no clear dominant orientation 3) VBN labels (derived from the profile and inferred values; interpreted for belief/accuracy judgments): – AC: Low | Medium | High. How likely does the profile indicate they have strong perceived severity of climate impacts? – AR: Low | Medium | High. How likely does the profile suggest they attribute climate change to human activity rather than natural causes? – PN: Low | Medium | High. How much internal pressure would the profile feel to align with pro-environmental positions? 4) synthesis: 2–3 sentences (max 50 words) citing which evidence supports the labels, and 2–3 sentences inferring how these VBN factors drive the final answer. Output JSON ONLY: {
"evidence": ["...", "..."],
"values": "self-enhancement | mixed | self-transcendence",
"VBN": {"AC":"Low|Medium|High","AR":"Low|Medium|High","PN":"Low|Medium|High"},
"synthesis": "..."
Stage 2 – System Same as baseline.
Stage 2 – User (belief as an example) # Profile:
Here is your profile: [PROFILE]
Here is a VBN plan extracted from the profile: __PLAN_JSON__ Answer the survey question below by applying the VBN plan. Use the VBN plan as a reference framework, but allow nuanced responses based on the specific question context. Follow the required response format EXACTLY. How accurate do you think these statements are? Answer on a scale from 0 to 100, where 0 means “not at all accurate” and 100 means “extremely accurate”. Questions: Q1: Taking action to fight climate change is necessary to avoid a global catastrophe. Q2: Human activities are causing climate change. Q3: Climate change poses a serious threat to humanity. Q4: Climate change is a global emergency. Respond in the following JSON format only, with no additional text: {"Q1": <number>, "Q2": <number>, "Q3": <number>, "Q4": <number>}
Figure S1: Intervention-level causal bias and cross-task correlation structure for Gemini. a. Signed ATE error for each of 11 interventions, grouped by intervention category (deductive reasoning, situational simulation, cultural and group norms) and outcome. b. Cross-task correlation matrices for human data and three prompting methods. Lower triangles show individual-level response correlations; upper triangles show intervention-level ATE correlations.
Figure S2: Intervention-level causal bias and cross-task correlation structure for Claude. a. Signed ATE error for each of 11 interventions, grouped by intervention category (deductive reasoning, situational simulation, cultural and group norms) and outcome. b. Cross-task correlation matrices for human data and three prompting methods. Lower triangles show individual-level response correlations; upper triangles show intervention-level ATE correlations.
Figure S3: Divergence of country-level descriptive and causal error and their inequality profiles for Gemini. a. Country-level scatterplots of baseline MAE against ATE error for each outcome. b. Countries ranked by MAE (top and bottom 10). c. Countries ranked by ATE error (top and bottom 10). d. Between-group differences in MAE by country characteristics. e. Between-group differences in ATE error by country characteristics. f. Between-group differences in MAE by demographic characteristics. g. Between-group differences in ATE error by demographic characteristics.
Figure S4: Divergence of country-level descriptive and causal error and their inequality profiles for Claude. a. Country-level scatterplots of baseline MAE against ATE error for each outcome. b. Countries ranked by MAE (top and bottom 10). c. Countries ranked by ATE error (top and bottom 10). d. Between-group differences in MAE by country characteristics. e. Between-group differences in ATE error by country characteristics. f. Between-group differences in MAE by demographic characteristics. g. Between-group differences in ATE error by demographic characteristics.